OpenStudy (anonymous):

simplify √108 -3 √50+5 √12+4√18

OpenStudy (campbell_st):

this can be rewritten as \[\sqrt{36 \times 3} - \sqrt{25 \times 2} + 5\sqrt{4 \times 3} + 4\sqrt{9 \times 2}\] which is equal to \[6\sqrt{3} - 5\sqrt{2} + 5\times2\sqrt{3} + 4\times3\sqrt{2}\] which is \[6\sqrt{3} - 5\sqrt{2} + 10\sqrt{3} + 12\sqrt{2}\] collect like terms for the answer

OpenStudy (anonymous):

\[16\sqrt{3}-12\sqrt{2}\]

OpenStudy (campbell_st):

no its \[16\sqrt{3} + 7\sqrt{2}\]

OpenStudy (campbell_st):

as its \[(12 - 5)\sqrt(2)\]

OpenStudy (anonymous):

its actually \[16\sqrt{3}-3\sqrt{2}\]

OpenStudy (anonymous):

my bad srry !!

OpenStudy (campbell_st):

oops missed a 3.. hope it helps